OpenStudy (anonymous):

Rewrite 2x + 3y = 9 in slope-intercept form and then graph the line.

OpenStudy (anonymous):

slope intercept form is: \[y=mx+b\] The problem is currently in standard form: \[Ax+By=C\] First, you must find the slope: \[m=\frac{ -A }{ B}\] For example: 3x - 4y = 12 \[m=\frac{ -3 }{ 4 } = \frac{ 3 }{ 4 }\] Subtract -3x from both sides, leaving -4y by itself: 3x - 4y = 12 -4y = 12 - 3x Divide -4 on all sides: \[y=-3+\frac{ 3 }{ 4 }x\]

OpenStudy (anonymous):

* Forgot to put a negative symbol for B, so this is wrong: \[m=\frac{ -3 }{ 4 }=\frac{ 3 }{ 4 }\] It's supposed to be: \[m=\frac{ -3 }{ -4 }=\frac{ 3 }{ 4 }\] Sorry for any confusions (:

OpenStudy (anonymous):

so is it still the same answer the y=-3+3/4

OpenStudy (anonymous):

Yes it is.